OpenStudy (anonymous):

You are contemplating investing $50,000 between two different investments. One of the investments, a stable one, returns 5% annually. A more risky one returns 9% annually. You need a return of at least 7.5%. what is the maximum amount of the money that you can invest at the 5% rate?

OpenStudy (anonymous):

@Loser66 Need some assistance if you have the time.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

I guess one way to solve this is let X = amount at 5%, P=$50,000, so (P-X) = amount invested at 9%. You want X(1.05) + (P-X)(1.09) = P(1.075) Simplify and solve for X, and I think that'll work.